OpenStudy (anonymous):

Solve using the quadratic formula: 5m^2+2m+2=0

hartnn (hartnn):

Compare your quadratic equation with \(am^2+bm+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{m_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

***roots of m